Find the length of the arc on the unit circle with the given central angles 315º


1
Expert's answer
2021-01-20T03:14:54-0500

"L=2\pi r*\frac{\theta}{360}"

"L=2\pi*1*\frac{315}{360}=\frac{7}{4}\pi=5.50."